The due date for a project is 20 weeks. If the expected activity times on a noncritical

path is 19 weeks, what is the probability that this path will take more than 20 weeks? The variance of

the activities on this path is 4.16.

The probability is ......... percent. (Enter the closest whole number)
Answer #1

z value (<=T) = (T - Expected time)/standard deviation

Expected time = 19

standard deviation = sqrt(Variance) = sqrt(4.16) = 2.04

z value (<=20) = (20-19)/2.04 = 0.49

Probability for completion in less than 20 weeks basis z value of 0.49 is 0.6880 or 68.80% (Using z table)

Probability that path will take more than 20 weeks is = 1 - 68.80% = 31.20% or 31%